40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 15 20:00:23 2020, Koji, Summary, General, HEPA AC cord replacement P_20201015_200732.jpgP_20201015_200752.jpgP_20201015_202615.jpgP_20201015_204234.jpg
    Reply  Thu Apr 1 23:55:33 2021, Koji, Summary, General, HEPA AC cord replacement 
       Reply  Fri Apr 2 15:17:23 2021, gautam, Summary, General, HEPA AC cord replacement 
          Reply  Tue Apr 6 21:17:04 2021, Koji, Summary, General, PSL HEPA investigation 20210406211741_IMG_0554.jpeg20210406211840_IMG_0555.jpeg20210406211850_IMG_0556.jpeg
Message ID: 15630     Entry time: Thu Oct 15 20:00:23 2020     Reply to this: 15989
Author: Koji 
Type: Summary 
Category: General 
Subject: HEPA AC cord replacement 

The AC cord from the PSL HEPA variac to the junction box was replaced.
Now the HEPA is running at 70%


Showed up at the 40m at 7pm

Preparation

  • Closed the PSL shutter.
  • Closed the innolight shutter
  • Turned off the HEPA mains switch
  • Checked the HEPA fan rating: 115V 4.5A.
  • Brought the thickest power cord from the wall stock: the rating is 125V 15A. This should sufficiently hold two HEPAs.

Cable Replacing

  • Rechecked the wire connection. The new cord has green/black/white wires. And the colors agree with the color of the wires in the junction box.
  • Removed the existing cord.
  • Attached the new cord.
  • Checked the variac AC plug. The terminals in the plug look normal and the AC plug looked sufficiently rigid.
  • Checked the connection again. = OK

Testing

  • Turned on the HEPA mains switch
  • VairAC turned to 70%
  • Checked the air flow - The HEPA fans are sucking the air = OK

Closing the work

  • Closed the junction box.
  • Cleaned up the roof
  • Opend the innolight shutter
  • Opened the PSL shutter
  • Locked the PMC
  • Locked the IMC  - found the transmission was ~80% of the pre-work due to misalignment of the PMC
  • Aligned the PMC - this recovered the IMC REFL of ~5.2 when the IMC was unlocked

Leaving the 40m at 9:30pm

Memo: 40m wiring/Mask/Camera/Red Pitaya/Particle Counter

Attachment 1: P_20201015_200732.jpg  1.932 MB  Uploaded Thu Oct 15 22:39:56 2020  | Hide | Hide all
P_20201015_200732.jpg
Attachment 2: P_20201015_200752.jpg  56 kB  Uploaded Thu Oct 15 22:42:10 2020  | Hide | Hide all
P_20201015_200752.jpg
Attachment 3: P_20201015_202615.jpg  120 kB  Uploaded Thu Oct 15 22:45:08 2020  | Hide | Hide all
P_20201015_202615.jpg
Attachment 4: P_20201015_204234.jpg  1.130 MB  Uploaded Thu Oct 15 22:46:51 2020  | Hide | Hide all
P_20201015_204234.jpg
ELOG V3.1.3-